OpenStudy (mtimko):

A rectangular cast-iron colum has a cross-sectional area of 25.0 m x 25.0 m. If the column supports a weight of 11,250 N, then how much stress is on the column (in Pascals)?

OpenStudy (michele_laino):

the unit stress is given by the subsequent formula: \[\Large \sigma = \frac{F}{S}\] where F= 11,250 newtons and S=25*25=...m^2

OpenStudy (michele_laino):

unitary* stress